OpenStudy (anonymous):

sin x = sin(x/2), evaluate x at range 0<=x<=2pi

OpenStudy (anonymous):

i first use the double angle identites then i get: 2sin(x/2)cos(x/2) = sin (x/2) then i change cos(x/2) to \[\pm \sqrt{1 - \sin^{2}(\frac{ x }{ 2 })}\] then i let y = sin (x/2), then i come to this expression \[y^{2}(3-4y^{2}) = 0, y=0, \pm (\frac{ \sqrt{3} }{ 2 })\] then my x = 0, 2pi, 2pi/3 and 4pi/3 but when i plug back x = 4pi / 3 in to the sin x = sin(x/2), i find that they re not equate, but instead they re in opposite sign but same magnitude. Anyone know why this approach is wrong?

myininaya (myininaya):

you know you could have made this easier...

myininaya (myininaya):

We have sin(x)=sin(x/2) Set the insides equal. Just to make sure we can find the most solutions. Let's use the fact that sin(x)=cos(pi/2-x) So cos(pi/2-x)=cos(pi/2-x/2) Set insides equal. Then let's use the fact cos is is even so cos(-x)=cos(x) So cos(x-pi/2) =cos(pi/2-x/2) set insides equal. So we could also look at cos(pi/2-x)=cos(x/2-pi/2) just make sure these x's are in the given interval now let me take a look at your process.

myininaya (myininaya):

oh wow i like the way you started

myininaya (myininaya):

\[2\sin(x/2)\cos(x/2)-\sin(x/2)=0\] now factor \[\sin(x/2)(2\cos(x/2)-1)=0\] set both factors equal to 0 \[\sin(x/2)=0 \] \[ 2\cos(x/2)-1=0\]

myininaya (myininaya):

so 0<=x<=2pi so 0/2<=x/2<=pi let x/2=u so we know that u is between 0 and pi including those endpoints. sin(u)=0 when u=? on the interval [0,pi] cos(u)=1/2 when u=? on the interval [0,pi] this way right here is way better that writing the equation like 4 different ways like I did. lol.

myininaya (myininaya):

than*

myininaya (myininaya):

You should get 3 solutions from this.

myininaya (myininaya):

Use the unit circle. Good state @hei .:)

OpenStudy (anonymous):

ty

myininaya (myininaya):

Hey but the reason your are going to have extra solutions (that are not actually solutions) is because you squared a radical.

myininaya (myininaya):

Just like if we had \[\sqrt{x+2}=-5\] square both sides x+2=25 x=23 this is not a solution

myininaya (myininaya):

\[\sqrt{x^2-4x+10}=5 \] square both sides \[x^2-4x+10=25\] \[x^2-4x-15=0 \] we get \[x=\frac{4 \pm \sqrt{16-4(-15)}}{2}=\frac{4 \pm \sqrt{16+60}}{2}\] But we have to plug back in to see if they both solutions because the square root functions had range [0, infinity) and not (-infinity, infinity)

myininaya (myininaya):

Does this make sense?

myininaya (myininaya):

Here is an easy problem to see what I mean: \[\sqrt{x-1}=x-7\] square both sides \[x-1=(x-7)^2 \] \[x-1=x^2-14x+49\] \[0=x^2-15x+50\] \[0=(x-5)(x-10) => x=5 \text{ or } x=10 \] But we have to check since we got rid of even root. \[\sqrt{10-1}=10-7 \text{ is true but } \sqrt{5-1} \neq 5-7 \]

myininaya (myininaya):

so the only solution is x=10.. So since you got rid of even root in your problem you must go back and check your solutions.

OpenStudy (anonymous):

I see, thanks again, a very good explanation